A daily production function for calculators is Q=12L^2 - L^3.

a) Write out an equation for the marginal product of labor.

b) Write out an equation for the average product of labor.
Answer #1

Q= 12 L2 - L3

(a) Marginal product of labor , MPL= dQ/dL = 12(2)L2-1 - (3)L3-1

= 24 L - 3L2

(b) Average product of labor , APL= Q/L = (12 L2 - L3)/L

= 12L - L2